How many possible sample sizes of 4 can be drawn from a population of size 10?

 


1
Expert's answer
2021-03-09T07:03:43-0500

"C^k_n=(^n_k)=\frac{n!}{k!(n-k)!}\\n=10, \ k=4\\\frac{10!}{4!(10-4)!}=210"

Answer: 210.